# AlignGPT: Multi-modal Large Language Models with Adaptive Alignment Capability

### v1

@@ 3.3 Adaptive Alignment-based Instruction-tuning
l|l|c|ccccccc
Method LLM Res. POPE MME MMB MMBCN SEEDI LLaVAW MM-Vet
BLIP-2 Vicuna-13B 224 85.3 1293.8 - - 46.4 38.1 22.4
InstructBLIP Vicuna-7B 224 - - 36.0 23.7 53.4 60.9 26.2
InstructBLIP Vicuna-13B 224 78.9 1212.8 - - - 58.2 25.6
Shikra Vicuna-13B 224 - - 58.8 - - - -
IDEFICS-9B LLaMA-7B 224 - - 48.2 25.2 - - -
IDEFICS-80B LLaMA-65B 224 - - 54.5 38.1 - - -
MiniGPT-v2 LLaMA2-7B 448 85.1 1332.1 43.1 29.1 52.3 - -
Qwen-VL Qwen-7B 448 - - 38.2 7.4 56.3 - -
Qwen-VL-Chat Qwen-7B 448 - 1487.5 60.6 56.7 58.2 - -
LLaVA-1.5 Vicuna-7B 336 85.9 1510.7 64.3 58.3 66.2 63.4 30.5
LLaVA-1.5 Vicuna-13B 336 85.9 1531.3 67.7 63.6 68.2 70.7 35.4
AlignGPT Vicuna-7B 336 86.0 1527.4 67.3 59.9 66.5 68.4 30.8
AlignGPT Vicuna-13B 336 86.2 1572.0 69.5 63.7 67.8 75.2 35.6
4 Experiments
4.1 Experimental Settings
Datasets.
For a fair comparison, we use the same pre-training and instruction dataset as the LLaVA-1.5 DBLP:journals/corr/abs-2310-03744 . It mainly includes 558K caption pairs for modality alignment and 665K single- or multi-round conversations for instruction-tuning. Besides, we evaluate AlignGPT on a range of academic visual question answering (VQA) tasks and recent benchmarks designed specifically for MLLMs. This evaluation spans 12 benchmarks, including VQAV2 DBLP:conf/cvpr/GoyalKSBP17 , GQA DBLP:conf/cvpr/Gurari0SGLGLB18 , VizWiz DBLP:conf/cvpr/Gurari0SGLGLB18 , SQAI (ScienceQA-IMG) DBLP:conf/nips/LuMX0CZTCK22 , TextVQA DBLP:conf/cvpr/SinghNSJCBPR19 , POPE DBLP:conf/emnlp/LiDZWZW23 , MME DBLP:journals/corr/abs-2306-13394 , MMB (MMBench), MMBCN (MMBench-Chinese) DBLP:journals/corr/abs-2307-06281 , SEEDI (SEED-Bench-IMG) DBLP:journals/corr/abs-2307-16125 , LLaVAW (LLaVA-Bench-in-the-Wild) DBLP:conf/nips/LiuLWL23a , and MM-Vet DBLP:journals/corr/abs-2308-02490 datasets.
Implementation Details.
We adopt a ViT dosovitskiy2021an model pre-trained with CLIP DBLP:conf/icml/RadfordKHRGASAM21 as a vision encoder to effectively process visual inputs. On the language side, Vicuna chiang2023vicuna is utilized to handle multimodal features, ensuring a cohesive integration of text and visual data. In the pre-training phase, both the visual backbone and the large language model of AlignGPT remain frozen, with only the parameters of the linear projection layer and alignment vectors being trained. During instruction-tuning phase, we freeze the linear projection layer, alignment vectors, and visual backbone, while adjusting the parameters of the large language model and the gate network. The global batch sizes for the two phases are set at 256 and 128 respectively, with DeepSpeed DBLP:conf/sc/RajbhandariRRH20 using ZeRO2 and ZeRO3 strategies accordingly. Regarding our training methodology, we conduct a single epoch of optimization for all models using the AdamW loshchilov2018decoupled optimizer coupled with a cosine learning schedule. Moreover, we initiate pre-training and instruction-tuning with learning rates of 1e-3 and 2e-5, respectively. The framework is trained on 8 A800 GPUs with 80GB memory.
4.2 Compared Methods
We chose a diverse set of representative multimodal large language models (MLLMs) as our baselines, including BLIP-2 DBLP:conf/icml/0008LSH23 , InstructBLIP DBLP:conf/nips/Dai0LTZW0FH23 , Shikra DBLP:journals/corr/abs-2306-15195 , IDEFICS DBLP:conf/nips/LaurenconSTBSLW23 , MiniGPT-v2 DBLP:journals/corr/abs-2310-09478 , Qwen-VL DBLP:journals/corr/abs-2308-12966 , Qwen-VL-Chat DBLP:journals/corr/abs-2308-12966 , and LLaVA-1.5 DBLP:journals/corr/abs-2310-03744 .
5 Results and Analysis
5.1 Main Results
Visual Question Answering.
We evaluate AlignGPT using five popular academic benchmarks, as detailed in Table 3.3. Despite using less training data, the AlignGPT-7B demonstrates competitive performance, surpassing other generalist models including InstructBLIP-13B, Shikra-13B, and IDEFICS-80B on most datasets, except for LLaVA-1.5-13B. These results verify the rationality of the structural design of our model. Moreover, considering that AlignGPT utilizes the same training dataset as LLaVA-1.5, it is evident that AlignGPT-7B outperforms LLaVA-1.5-7B across all evaluation datasets, and AlignGPT-13B also surpasses LLaVA-1.5-13B on the majority of datasets. This demonstrates that our approach has effectively enhance the alignment capabilities of multimodal large language models. The fly in the ointment is that AlignGPT-13B does not perform as well as Qwen-VL on the TextVQA dataset. This may stem from the fact that TextVQA is a text-centric QA task, as it requires identifying text in images to answer questions. AlignGPT is tailored to boost multimodal alignment and might not exhibit strong results in text-focused scenarios.
MLLM-oriented Multi-modal Benchmarks.
We apply AlignGPT to seven recent popular multimodal benchmarks, as shown in Table 3.3. We discover that, apart from LLaVA-1.5-13B, AlignGPT-7B surpassed all previous multimodal models. This shows that our model has strong generalization ability. Additionally, compared to LLaVA-1.5-13B, AlignGPT-13B has shown improvements on most datasets, particularly achieving good advancements on the MME, MMB, and LLaVAW benchamrks. This further validates the efficacy of both global and local alignment capabilities.
Table 3: Influence of different number of alignment level.
{tabu}
l|c|ccccccccc Method Alignment Level VQAv2 GQA VisWiz SQAI TextVQA POPE MME MMB SEEDI
@@ 5.1 Main Results
Table 4: Influence of local and global alignment.
@@
c|ccc|cccccccccc Settings Average Local Global VQAv2 GQA VisWiz SQAI TextVQA POPE MME MMB SEEDI
@@
5.2 Ablation Study
Without loss of generality, we choose AlignGPT-7b for the ablation study to investigate the effects of different components.
Impact of Number of Alignment Levels.
To investigate the effect of the number of alignment levels $N$ on AlignGPT, we vary the value of $N$ in the range of [4, 10] with a step size of 2. Table  5.1 shows the performance of AlignGPT with different $N$ on nine datasets. Actually, AlignGPT can achieve good results at $N=4$, and their performance remains stable as the number of alignment levels increases. Depending on the trajectory of the curve, their performance has an initial upward trend and then flattens out. These observations indicate that AlignGPT can improve the alignment capabilities of multi-modal large language models based on a small number of alignment levels. Finally, according to the trend of the curve, we set $N$ to 8.
Impact of Local and Global Alignment.
During the instruction-tuning phase, we assign global alignment capabilities and local alignment capabilities to the instructions of each task. Here, we explore the role of separate global alignment capabilities and local alignment capabilities on AlignGPT. Among them, “Local” refers to the local alignment capabilities derived by assigning different weights to various local alignment vectors using a gate network. “Global” denotes the global alignment capabilities, and “Average” represents the local alignment capabilities obtained by assigning equal weights to each local alignment vector. The performance of these four strategies (settings a-d) is presented in Table 5.1. As we can see, setting(a) and setting(b) demonstrate divergent performances in downstream tasks, which can be attributed to the different demands these tasks place on global and local alignment capabilities. It is worth noting that setting (a) and setting (b) perform worse than our final approach (setting d) on most datasets, which verifies the necessity and superiority of the combination of global and local alignment capabilities. Additionally, the performance of setting (c) is inferior to that of setting (d), a possible reason being that the demands for local alignment capabilities in each downstream task are dynamically variable.
Table 5: Influence of different input image resolutions.
@@ 5.2 Ablation Study
l|l|c|ccccccc Method LLM Resolution VQAv2 GQA SQAI TextVQA POPE MMB SEEDI
@@
{tabu}
l|l|c|ccccccc Method LLM Resolution VQAv2 GQA SQAI MME MMB MMBCN SEEDI
AlignGPT LLaMA2-7B-Chat 336 79.1 62.9 65.9 1500.8 66.6 57.9 66.4
AlignGPT Vicuna-7B 336 79.1 62.9 68.5 1527.4 67.3 59.9 66.5
AlignGPT LLaMA3-8B-Base 336 79.6 63.1 70.4 1539.7 72.0 67.7 68.2
5.3 Discussion
Impact of different input image resolutions.
Image resolution plays a crucial role in vision-language tasks as higher resolutions help reduce image blurring and enhance the understanding of image-text alignment. To evaluate the impact of resolution changes on the performance of multimodal tasks, we increase the image resolution from 336 to 1008, with the resulting performance changes detailed in Table 5.2. The study results show that higher image resolutions can improve model performance on most multimodal tasks. For example, the score for VQAv2 increased from 79.1 to 79.8, while the score for TextVQA rose from 58.4 to 60.3. Meanwhile, the performance of the POPE improve by 0.8. These results highlight that appropriately increasing image resolution is an effective strategy for enhancing performance in studies of multimodal large language models.
Impact of different large language models.
We also explore the impact of the large language model on the performance of AlignGPT, specifically testing three models: LLaMA-2-7B-Chat, Vicuna-v1.5-7B, and the latest LLaMA-3-8B-Base. The results are shown in Table 5.2. Initially, we observe that LLaMA-3-8B-Base achieves the best performance, followed by Vicuna-v1.5-7B, with LLaMA-2-7B-Chat performing the worst, which is reasonable given LLaMA-3-8B-Base’s larger parameter size and richer training data. Besides, we find that compared to the VQA tasks, Vicuna-v1.5-7B performs better on multimodal benchmarks such as MME, MMB, and SEEDI than LLaMA-2-7B-Chat, possibly because Vicuna-v1.5-7B underwent supervised instruction-tuning with ShareGPT data, and ShareGPT contains some background knowledge related to downstream tasks.
5.4 Qualitative Results
Figure 3 presents a comparative analysis of our model with MiniGPT-v2 DBLP:journals/corr/abs-2310-09478 and LLaVA-1.5 DBLP:journals/corr/abs-2310-03744 . When a user submits an image alongside the instruction “Can you see the bright blue feathers on the bird’s wing?”, MiniGPT-v2 and LLaVA-1.5 both return an incorrect answer “Yes”. In contrast, our model produces accurate result “No”, thereby demonstrating that AlignGPT can effectively enhance the model’s alignment capability. In Figure 4, we further demonstrate the responses of AlignGPT under different levels of alignment capability. We find that with lower alignment levels, the model may only focus on certain regions of the image, resulting in an undercount of the total number of apples; whereas with higher alignment levels, the model considers the entire image area, thus achieving accurate apple quantity estimation. This finding once again underscores the necessity of enhancing the alignment capability of MLLMs.
@@ 6 Conclusion
In this paper, we propose AlignGPT, a novel multimodal large language model designed to bolster the alignment capabilities of MLLMs. Our approach involves utilizing the alignment level of data as a control signal during pre-training to effectively handle the varying degrees of alignment in image-text pairs. Subsequently, in the instruction-tuning phase, we begin by exploiting these control signals to shape different levels of alignment capabilities. Continuing from this, we go beyond assigning global alignment capabilities to instructions of each task; we also dynamically configure distinct local alignment capabilities based on the specific demands of each instruction. Results from numerous experiments indicate that our AlignGPT achieves better performance than other state-of-the-art MLLMs. Further analysis validates the superiority of our AlignGPT.

### v2

@@ 3.4 Adaptive Alignment-based Instruction-tuning
l—c—ccccccccc Method Alignment Level VQAv2 GQA VisWiz SQAI TextVQA POPE MME MMB SEEDI
@@
Table 3: Influence of different number of alignment level.
@@
c—ccc—cccccccccc Settings Average Local Global VQAv2 GQA VisWiz SQAI TextVQA POPE MME MMB SEEDI
@@
Table 4: Influence of local and global alignment.
4 Experiments
4.1 Experimental Settings
Datasets.
For a fair comparison, we use the same pre-training and instruction dataset as the LLaVA-1.5 [24]. It mainly includes 558K caption pairs for modality alignment and 665K single- or multi-round conversations for instruction-tuning. Besides, we evaluate AlignGPT on a range of academic visual question answering (VQA) tasks and recent benchmarks designed specifically for MLLMs. This evaluation spans 12 benchmarks, including VQAV2 [16], GQA [18], VizWiz [18], SQAI (ScienceQA-IMG) [28], TextVQA [37], POPE [23], MME [15], MMB (MMBench), MMBCN (MMBench-Chinese) [26], SEEDI (SEED-Bench-IMG) [21], LLaVAW (LLaVA-Bench-in-the-Wild) [25], and MM-Vet [44] datasets.
Implementation Details.
We adopt a ViT [12] model pre-trained with CLIP [34] as a vision encoder to process visual inputs. On the language side, Vicuna [9] is utilized to handle multimodal features, ensuring a cohesive integration of text and visual data. In the pre-training phase, both the visual backbone and the large language model of AlignGPT remain frozen, with only the parameters of the linear projection layer and alignment embeddings being trained. During instruction-tuning phase, we freeze the alignment embeddings and visual backbone, while adjusting the parameters of the linear projection layer, large language model, and the gate network. The global batch sizes for the two phases are set at 256 and 128 respectively, with DeepSpeed [35] using ZeRO2 and ZeRO3 strategies accordingly. Regarding our training methodology, we conduct a single epoch of optimization for all models using the AdamW [27] optimizer coupled with a cosine learning schedule. Moreover, we initiate pre-training and instruction-tuning with learning rates of 1e-3 and 2e-5, respectively. The framework is trained on 8 A800 GPUs with 80GB memory.
4.2 Compared Methods
We chose a diverse set of representative MLLMs as our baselines, including BLIP-2 [22], InstructBLIP [10], Shikra [6], IDEFICS [20], MiniGPT-v2 [5], Qwen-VL [3], Qwen-VL-Chat [3], and LLaVA-1.5 [24].
@@ 4.2 Compared Methods
l—l—c—ccccccc Method LLM Resolution VQAv2 GQA SQAI TextVQA POPE MMB SEEDI
@@
Table 5: Influence of different input image resolutions.
{tabu}
l—l—c—ccccccc Method LLM Resolution VQAv2 GQA SQAI MME MMB MMBCN SEEDI
AlignGPT LLaMA2-7B-Chat 336 79.1 62.9 65.9 1500.8 66.6 57.9 66.4
AlignGPT Vicuna-v1.5-7B 336 79.1 62.9 68.5 1527.4 67.3 59.9 66.5
AlignGPT LLaMA3-8B-Base 336 79.6 63.1 70.4 1539.7 72.0 67.7 68.2
@@
5 Results and Analysis
5.1 Main Results
MLLM-oriented Multi-modal Benchmarks.
We apply AlignGPT to seven recent popular multimodal benchmarks, as shown in  Sec. 3.4. We discover that, apart from LLaVA-1.5-13B, AlignGPT-7B surpassed all previous multimodal models. This shows that our model has strong generalization ability. Additionally, compared to LLaVA-1.5-13B, AlignGPT-13B has shown improvements on most datasets, particularly achieving good advancements on the MME, MMB, and LLaVAW benchamrks. This further validates the efficacy of both global and local alignment capabilities.
Visual Question Answering.
We evaluate AlignGPT using five popular academic benchmarks, as detailed in  Sec. 3.4. Despite using less training data, the AlignGPT-7B demonstrates competitive performance, surpassing other generalist models including InstructBLIP-13B, Shikra-13B, and IDEFICS-80B on most datasets, except for LLaVA-1.5-13B. These results verify the rationality of the structural design of our model. Moreover, considering that AlignGPT utilizes the same training dataset as LLaVA-1.5, it is evident that AlignGPT-7B outperforms LLaVA-1.5-7B across all evaluation datasets, and AlignGPT-13B also surpasses LLaVA-1.5-13B on the majority of datasets. This demonstrates that our approach effectively enhances the alignment capabilities of multimodal large language models. The fly in the ointment is that AlignGPT-13B does not perform as well as Qwen-VL on the TextVQA dataset. This may stem from the fact that TextVQA is a text-centric QA task, as it requires identifying text in images to answer questions. AlignGPT is tailored to boost multimodal alignment and might not exhibit strong results in text-focused scenarios.
Figure 5: The performance comparison of AlignGPT (random) and AlignGPT in downstream datasets.
@@ 5.1 Main Results
5.2 Ablation Study
Without loss of generality, we choose AlignGPT-7b for the ablation study to analyze the impact of various components.
Impact of Alignment Embedding Table.
We design an experiment to validate the role of the alignment embedding table. In the fine-tuning phase, we use a randomly initialized alignment embedding table instead of the embedding table obtained during the pre-training phase. Since both methods have the same number of parameters, we can clearly assess whether the parameter number of the alignment embedding table influences the improvement in model performance. The experimental results are shown in  Fig. 5. We find that the model using the randomly initialized alignment embedding table (referred to as AlignGPT (random)) shows a performance gap compared to AlignGPT. These results further confirm that the alignment information learned during pre-training is indeed the key factor in enhancing model performance, rather than the parameter number.
Impact of Number of Alignment Levels.
To investigate the effect of the number of alignment levels $N$ on AlignGPT, we vary the value of $N$ in the range of [4, 10] with a step size of 2.  Sec. 3.4 shows the performance of AlignGPT with different $N$ on nine datasets. Actually, AlignGPT can achieve good results at $N=4$, and their performance remains stable as the number of alignment levels increases. Depending on the trajectory of the curve, their performance has an initial upward trend and then flattens out. These observations indicate that AlignGPT can improve the alignment capabilities of multi-modal large language models based on a small number of alignment levels. Finally, according to the trend of the curve, we set $N$ to 8.
Impact of Local and Global Alignment.
During the instruction-tuning phase, we assign global and local alignment capabilities to the instructions of each task. Among them, “Local” refers to the local alignment capabilities derived by assigning different weights to various local alignment embeddings using a gate network. “Global” denotes the global alignment capabilities, and “Average” represents the local alignment capabilities obtained by assigning equal weights to each local alignment embedding. The performance of these four strategies (settings a-d) is presented in  Sec. 3.4. As we can see, setting (a) and setting (b) demonstrate divergent performances in downstream tasks, which can be attributed to the different demands these tasks place on global and local alignment capabilities. It is worth noting that setting (a) and setting (b) perform worse than our final approach (setting d) on most datasets, which verifies the necessity of the combination of global and local alignment capabilities. Moreover, the performance of setting (c) is inferior to that of setting (d), which may be due to the dynamically changing demands for local alignment capabilities across different downstream tasks.
5.3 Discussion
Impact of different input image resolutions.
Image resolution plays a crucial role in vision-language tasks as higher resolutions help reduce image blurring and enhance the understanding of image-text alignment. To evaluate the impact of resolution changes on the performance of multimodal tasks, we increase the image resolution from 336 to 1008, with the resulting performance changes detailed in  Sec. 4.2. The study results show that higher image resolutions can improve model performance on most multimodal tasks. For example, the score for VQAv2 increased from 79.1 to 79.8, while the score for TextVQA rose from 58.4 to 60.3. Meanwhile, the performance of the POPE improve by 0.8. These results highlight that appropriately increasing image resolution is an effective strategy for enhancing performance in studies of multimodal large language models.
Impact of different large language models.
We also explore the impact of the large language model on the performance of AlignGPT, specifically testing three models: LLaMA-2-7B-Chat, Vicuna-v1.5-7B, and the latest LLaMA-3-8B-Base. The results are shown in  Sec. 4.2. Initially, we observe that LLaMA-3-8B-Base achieves the best performance, followed by Vicuna-v1.5-7B, with LLaMA-2-7B-Chat performing the worst, which is reasonable given LLaMA-3-8B-Base’s larger parameter size and richer training data. Besides, we observe that Vicuna-v1.5-7B achieves superior performance over LLaMA-2-7B-Chat on multimodal benchmarks such as MME, MMB, and SEEDI, while showing comparable results on VQA tasks. This advantage might be due to Vicuna-v1.5-7B undergoing supervised instruction-tuning with ShareGPT data, which contains background knowledge relevant to downstream tasks.
5.4 Qualitative Results
Figure 6 presents a comparative analysis of our model with MiniGPT-v2 [5] and LLaVA-1.5 [24]. When a user submits an image alongside the instruction “Can you see the bright blue feathers on the bird’s wing?”, MiniGPT-v2 and LLaVA-1.5 both return an incorrect answer “Yes”. In contrast, our model produces accurate result “No”, thereby demonstrating that AlignGPT can effectively enhance the model’s alignment capability. In Figure 7, we further demonstrate the responses of AlignGPT under different levels of alignment capability. We find that with lower alignment levels, the model may only focus on certain regions of the image, resulting in an undercount of the total number of apples; whereas with higher alignment levels, the model considers the entire image area, thus achieving accurate apple quantity estimation. This finding once again underscores the necessity of enhancing the alignment capability of MLLMs.
@@ 6 Conclusion
In this paper, we propose AlignGPT, a novel multimodal large language model designed to bolster the alignment capabilities of MLLMs. Our approach involves utilizing the alignment level of data as a control signal during pre-training to effectively handle the varying degrees of alignment in image-text pairs. Subsequently, in the instruction-tuning phase, we begin by exploiting these control signals to shape different levels of alignment capabilities. Continuing from this, we go beyond assigning global alignment capabilities to instructions of each task; we also dynamically configure distinct local alignment capabilities based on the specific demands of each instruction. Results from numerous experiments indicate that our AlignGPT achieves better performance than other state-of-the-art MLLMs.
Limitations
The current study has two limitations: (1) This paper involves two modalities, i.e., text and image, while achieving AGI should also encompass video and audio, which requires us to do further research and exploration; (2) We propose a new perspective to enhance the alignment capability of MLLMs. However, there may be other methods to achieve this goal, which merit consideration in the future.